A simple pendulum is found to vibrate 50 times within 200s. When 1.5 m of its


length is reduced to certain length, it vibrates 50 times in 175 s. Find the


original length of the pendulum.

1
Expert's answer
2021-11-14T17:12:15-0500

The period of the original pendulum:


"T_1/N=2\pi\sqrt{L/g},\\\space\\
T_2/N=2\pi\sqrt{(L-l)/g}.\\\space\\
\frac{T_1}{T_2}=\sqrt{\frac L{L-l}},\\\space\\
\bigg(\frac{T_1}{T_2}\bigg)^2=\frac L{L-l},\\\space\\
L=l·\frac{T_1^2/T_2^2}{T_1^2/T_2^2-1}=6.4\text{ m}."